A machine costing $214,600 with a four-year life and an estimated $17,000 salvage value is installed in Luther Companys factory on January 1. The factory manager estimates the machine will produce 494,000 units of product during its life. It actually produces the following units: 123,000 in 1st year, 122,600 in 2nd year, 121,000 in 3rd year, 137,400 in 4th year. The total number of units produced by the end of year 4 exceeds the original estimate-this difference was not predicted. (The machine must not be depreciated below its estimated salvage value.) Required: Compute depreciation for each year (and total depreciation of all years combined) for the machine under each depreciation method. (Round your per unit depreciation to 2 decimal places. Round your answers to the nearest whole dollar.) Complete this question by entering your answers in the tabs below Straight Line Production DDB Compute depreciation for each year (and total depreciation of all years combined) for the machine under each Straight-line depreciation Straight-Line Depreciation Year Depreciation Expense Total
Units of Production Depreciable DepreciationDepreciation per unit Year Units Expense 2 4 Total
End of Period DDB Depreciation for the Period Depreciation Depreciation Accumulated Expense Depreciation Beginning of Book Value Year Period Book Rate Value 4
Answer #1
Straight Line depreciation :
year Depreciation Expense
1 49400 (214600-17000)/4
2 49400
3 49400
4 49400
Total 197600
Units of Production method of depreciation:
Year Depreciable units Depreciation per unit Depreciation expense
1 123000 (Cost - salvage)/Life units produced 49200
2 122600 (214600-17000)/494000 49040
3 121000 0.4 48400
4 137400 50960 (197600-49200-49040-48400)
Total 504000 197600
salvage 17000
Machine cost 214600
DDB depreciation method:
year Begin BV Depreciation rate Depreciation Exp. Acc. Dep. Book Value
0 214600
1 214600 SL dep. Rate * 2 = 100/4 = 25% *2 = 50% 107300 107300 107300
2 107300 50% 53650 160950 53650
3 53650 50% 26825 187775 26825
4 26825 26825-17000 = 9825 197600 17000 salvage
Total 197600